# On the Small Jumps of Lévy Processes and the Multivariate Dickman Distribution

### v1

@@ Abstract
We develop a Dickman approximation to the small jumps of multivariate Lévy processes and related stochastic integral processes. Further, we show that the multivariate Dickman distribution is the unique distribution satisfying a certain stability property and that a related stochastic process is self-similar. Along the way we develop applications to non-Gaussian OU-Processes, to the class of generalized multivariate gamma distributions, and to the popular Barndorff-Nielsen and Shephard stochastic volatility model.
@@ 1 Introduction
Lévy and related infinitely divisible processes are used in many application areas. When simulating a Lévy process, the difficulty typically lies in the simulation of the small jumps, as the large jumps follow a compound Poisson process, which is often easier to simulate. This has led to the study of approximations of small jumps by simpler processes. In the univariate case, an approximation of small jumps by Brownian motion was developed in [1]. This was extended to the multivariate case in [8], see also the discussion of applications to finance in Chapter 6 of [9]. For important situations where the Brownian motion approximation fails, [10] developed an approximation by a Dickman Lévy process in the univariate case. Recently, a multivariate generalization of the Dickman distribution was introduced in [6] and [15].
The goal of the current paper is three-fold. First, we develop a Dickman approximation to the small jumps of multivariate Lévy processes. Second, we extend these results to stochastic integral processes. Third, we show that the multivariate Dickman distribution is the unique distribution satisfying a certain stability property related to the stability of its jumps under scaling and that a related stochastic process is self-similar. To the best of our knowledge, the second and third of these were previously unknown even in the univariate case. Along the way we develop applications to non-Gaussian processes of Ornstein-Uhlenbeck-type (OU-processes), to the class of generalized multivariate gamma distributions, and to the popular Barndorff-Nielsen and Shephard (BNS) stochastic volatility model introduced in [5].
The rest of this paper is organized as follows. In Section 2 we review infinitely divisible distributions, Lévy processes, and stochastic integral processes and we prove a transfer theorem that allows one to transfer convergence results about sequences of Lévy processes to sequences of stochastic integral processes. This will be needed to prove our main results and may be of independent interest. In Section 3 we recall basic facts about Dickman and multivariate Dickman distributions and introduce the more general class of $\epsilon$-multivariate Dickman distributions. We then show that these are the only distributions that satisfy a certain stability property and that a related stochastic process is self-similar. In Section 4 we give our main results for approximating the small jumps of certain multivariate Lévy processes by Dickman Lévy processes and we extend these results to stochastic integral processes. In Section 5 we give an application to the simulation of generalized multivariate gamma distributions. In Section 6 we give an application to the BNS stochastic volatility model. Proofs are postponed to Section 7.
@@
2 Infinitely Divisible Distributions, Lévy Processes, Stochastic Integrals, and a Transfer Theorem
In which we review basic properties of infinitely divisible distributions and their associated Lévy processes, discuss stochastic integral processes, and prove a transfer theorem that allows one to transfer convergence results about sequences of Lévy processes to sequences of stochastic integral processes. The transfer theorem is needed to prove our main results and may be of independent interest.
$C_{\mu}(z)=-\langle z,Az\rangle+i\left\langle b,z\right\rangle+\int_{\mathbb{R}^{d}}\left(e^{i\left\langle z,x\right\rangle}-1-i\left\langle z,x\right\rangle 1_{[|x|\leq 1]}\right)M(\mathrm{d}x),\ \ \ z\in\mathbb{R}^{d},$ (1)
$A$ is a $d\times d$-dimensional covariance matrix called the Gaussian part, $b\in\mathbb{R}^{d}$ is the shift, and $M$ is the Lévy measure, which is a Borel measure on $\mathbb{R}^{d}$ satisfying
$M(\{0\})=0\text{ and }\int_{\mathbb{R}^{d}}(|x|^{2}\wedge 1)M(\mathrm{d}x)<\infty.$
The parameters $A$, $M$, and $b$ uniquely determine this distribution and we write $\mu=\mathrm{ID}(A,M,b)$. We call $C_{\mu}$ the cumulant generating function (cgf) of $\mu$. The normal distribution corresponds to the case where $M=0$ and is denoted $N(b,A)$.
Associated with every infinitely divisible distribution $\mu$ is a Lévy process $\{X_{t}:t\geq 0\}$, where $X_{1}\sim\mu$. Lévy processes are characterized by independent and stationary increments, càdlàg paths, stochastic continuity, and the initial condition $X_{0}=0$ with probability $1$. Standard Brownian motion is the Lévy process $\{W_{t}:t\geq 0\}$ with $W_{1}\sim N(0,I_{d})$, where $I_{d}$ is the $d\times d$-dimensional identity matrix. The jumps of a Lévy process are governed by its Lévy measure. Specifically, if $\{X_{t}:t\geq 0\}$ is a Lévy process with $X_{1}\sim\mathrm{ID}(A,M,b)$, then, for any $B\in\mathfrak{B}(\mathbb{R}^{d})$, $M(B)$ is the expected number of jumps that the process has in the time interval $[0,1]$ that fall inside set $B$.
@@ A Lévy process has finite variation if and only if $A=0$ and $M$ satisfies the additional condition
$\int_{\mathbb{R}^{d}}(|x|\wedge 1)M(\mathrm{d}x)<\infty.$ (2)
Through a slight abuse of terminology, we also say that the associated distribution $\mu$ has finite variation. In this case, the cgf can be written in the form
@@
When discussing convergence in distribution of infinitely divisible distributions and Lévy processes, the concept of vague convergence is fundamental. A portmanteau theorem giving several statements that are equivalent to vague convergence can be found in, e.g., [2]. The definition is as follows.
@@
Let $M_{0},M_{1},M_{2},\dots$ be a sequence of Lévy measures on $\mathbb{R}^{d}$. We say that $M_{n}$ converges vaguely to $M_{0}$ and write $M_{n}\stackrel{{\scriptstyle v}}{{\rightarrow}}M_{0}$ as $n\to\infty$ if $\lim\limits_{n\to\infty}\int_{\mathbb{R}^{d}}f(x)M_{n}(\mathrm{d}x)=\int_{\mathbb{R}^{d}}f(x)M_{0}(\mathrm{d}x)$, for every $f:\mathbb{R}^{d}\mapsto\mathbb{R}$ that is bounded, continuous, and vanishing on a neighborhood of 00.
We now turn to stochastic integral processes. Let ×$\mathbb{R}^{\ell\times d}$ be the space ×$\ell\times d$-dimensional matrices. Let $F(\cdot,\cdot):[0,\infty)^{2}\mapsto\mathbb{R}^{\ell\times d}$ such that $F(\cdot,t)$ is a Borel function for each $t\geq 0$. We write $F_{i,j}(\cdot,\cdot)$ to denote the $i,j$th component of $F$. Assume that, for $t\geq 0$ and each $i,j$,
$\int_{0}^{\infty}|F_{i,j}(s,t)|\vee|F_{i,j}(s,t)|^{2}\mathrm{d}s<\infty.$ (4)
In this case we say that $F$ is an integrable kernel on ×$\mathbb{R}^{\ell\times d}$. Let $Z=\{Z_{t}:t\geq 0\}$ be a $d$-dimensional Lévy process with $Z_{1}\sim\mu=\mathrm{ID}(A,M,b)$ and set
$X_{t}=\int_{0}^{\infty}F(s,t)\mathrm{d}Z_{s},$
where the stochastic integral is defined in the sense of Definition 2.20 in [28]. Many properties of the stochastic integral are given in [28]. We call $X=\{X_{t}:t\geq 0\}$ a stochastic integral process with background driving Lévy process (BDLP) $Z$ and integrable kernel $F$. For each $t\geq 0$, the random variable $X_{t}$ takes values in $\mathbb{R}^{\ell}$. Applying Proposition 2.7 in [28] and dominated convergence, shows that the characteristic function of $X_{t}$ is given by
$\mathrm{E}\left[e^{i\langle z,X_{t}\rangle}\right]=\exp\left\{\int_{0}^{\infty}C_{\mu}(F^{\top}(s,t)z)\mathrm{d}s\right\},\ \ z\in\mathbb{R}^{\ell},$ (5)
where $C_{\mu}$ is the cgf of $Z_{1}$. It follows that $X_{t}\sim\mathrm{ID}(A_{t},M_{t},b_{t})$, where $A_{t}=\int_{0}^{\infty}F(s,t)AF^{\top}(s,t)\mathrm{d}s$,
$M_{t}(B)=\int_{0}^{\infty}\int_{\mathbb{R}^{d}}1_{B\setminus\{0\}}(F(s,t)x)M(\mathrm{d}x)\mathrm{d}s,\ \ B\in\mathfrak{B}(\mathbb{R}^{d}),$
and
$b_{t}=\int_{0}^{\infty}\left(F(s,t)b+\int_{\mathbb{R}^{d}}F(s,t)x\left(1_{[|F(s,t)x|\leq 1]}-1_{[|x|\leq 1]}\right)M(\mathrm{d}x)\right)\mathrm{d}s.$
A common situation is when $\ell=d$ and $F(s,t)=f(s,t)I_{d}$, where $I_{d}$ is the identity matrix in $\mathbb{R}^{d\times d}$ and $f:[0,\infty)^{2}\mapsto\mathbb{R}$ is such that $f(\cdot,t)$ is a Borel function for each $t\geq 0$ satisfying appropriate integrability conditions. In this case we often write
$X_{t}=\int_{0}^{\infty}f(s,t)\mathrm{d}Z_{s}$
instead of $\int_{0}^{\infty}F(s,t)\mathrm{d}Z_{s}$. Perhaps the most famous stochastic integral process is when $f(s,t)=e^{-c(t-s)}1_{[0\leq s\leq t]}$ for some $c>0$. This corresponds to an OU-process, see [9] or [24], and the references therein for details. Another important, although somewhat trivial, example is when $f(s,t)=1_{[0\leq s\leq t]}$. Here the stochastic integral process is the same as the original BDLP.
We now give our transfer theorem. It shows continuity of the stochastic integral process with respect to the BDLP. We have not seen a result of this type in the literature, although related results can be found in [23], [29], Proposition 8.2 in [4], and Proposition 3.7 in [28].
Theorem 1.
For $n=0,1,2,\dots$ let $\{Z_{t}^{(n)}:t\geq 0\}$ be a sequence of Lévy processes on $\mathbb{R}^{d}$ with $Z_{1}^{(n)}\sim\mu^{(n)}=\mathrm{ID}(A^{(n)},M^{(n)},b^{(n)})$ and let $X^{(n)}_{t}=\int_{0}^{\infty}F(s,t)\mathrm{d}Z_{s}^{(n)}$ for $t\geq 0$, where $F$ is an integrable kernel on ×$\mathbb{R}^{\ell\times d}$. Assume that there is a sequence of real numbers $\{c_{n}\}$ with $c_{n}Z_{1}^{(n)}\stackrel{{\scriptstyle d}}{{\rightarrow}}Z_{1}^{(0)}$. If either:
(i) there exists an $h\geq 0$ with $M_{h}:=\limsup_{n\to\infty}|c_{n}|\int_{|c_{n}x|>h}|x|M^{(n)}(\mathrm{d}x)<\infty$, or
(ii) there exist a function $a:[0,\infty)\mapsto[0,\infty)$ with $F(s,t)=0$ for all $s>a(t)$,
then $\{c_{n}X_{t}^{(n)}:t\geq 0\}\stackrel{{\scriptstyle fdd}}{{\rightarrow}}\{X_{t}^{(0)}:t\geq 0\}$.
Note that (ii) always holds when $X_{t}=\int_{0}^{t}F(s,t)\mathrm{d}Z_{s}=\int_{0}^{\infty}1_{[0\leq s\leq t]}F(s,t)\mathrm{d}Z_{s}$. In particular, it always holds for OU-processes.
3 $\epsilon$-Multivariate Dickman Distributions, Stability, and Self-Similarity
In which Dickman and multivariate Dickman distributions are reviewed, the more general class of $\epsilon$-multivariate Dickman distributions is introduced and shown to be characterized by a certain stability property, and a related process is shown to be self-similar.
@@ 3 $\epsilon$-Multivariate Dickman Distributions, Stability, and Self-Similarity
A random variable $X$ on $\mathbb{R}$ is said to have a generalized Dickman distribution if
$X\stackrel{{\scriptstyle d}}{{=}}U^{1/\theta}(X+1),$
@@
where $\theta>0$ and $U\sim U(0,1)$ is independent of $X$ on the right side. We denote this distribution by $\mathrm{GD}(\theta)$. When $\theta=1$, it is just called the Dickman distribution. Many properties of this distribution are discussed in the surveys [21], [20], and [14]. A multivariate generalization of the Dickman distribution was recently introduced in [6]. It was further studied in [15], where many properties were derived and several approaches for simulation were studied. A random variable $X$ on $\mathbb{R}^{d}$ is said to have a multivariate Dickman (MD) distribution if
$X\stackrel{{\scriptstyle d}}{{=}}U^{1/\theta}(X+\xi),$ (6)
where $\theta>0$ and $X,\xi,U$ are independent on the right side with $U\sim U(0,1)$ and $\xi\sim\sigma_{1}$ for some probability distribution $\sigma_{1}$ on $\mathbb{S}^{d-1}$. Let $\sigma=\theta\sigma_{1}$ and note that $\theta=\sigma(\mathbb{S}^{d-1})$ and $\sigma_{1}=\sigma/\sigma(\mathbb{S}^{d-1})$. Thus, there is no loss of information when working with $\sigma$ instead of $\theta$ and $\sigma_{1}$. We write $\mathrm{MD}(\sigma)$ to denote this distribution. The generalized Dickman distribution $\mathrm{GD}(\theta)$ corresponds to the multivariate Dickman distribution $\mathrm{MD}(\sigma)$ when the dimension $d=1$ and $\sigma=\theta\delta_{1}$. In [15] it is shown that the multivariate Dickman distribution is infinitely divisible. Before giving its Lévy measure, we introduce the more general class of $\epsilon$-multivariate Dickman distributions.
@@
For $\epsilon>0$, an infinitely divisible distribution $\mu$ on $\mathbb{R}^{d}$ is said to be an $\epsilon$-multivariate Dickman distribution if $\mu=\mathrm{ID}_{0}(D^{\epsilon},0)$, where
@@
$D^{\epsilon}(B)=\int_{\mathbb{S}^{d-1}}\int_{0}^{\epsilon}1_{B}(rs)r^{-1}\mathrm{d}r\sigma(\mathrm{d}s),\ \ \ B\in\mathfrak{B}(\mathbb{R}^{d})$ (7)
for some finite Borel measure $\sigma$ on $\mathbb{S}^{d-1}$. We denote this distribution by $\mathrm{MD}^{\epsilon}(\sigma)$ and we refer to $\sigma$ as the spectral measure. The multivariate Dickman distribution corresponds to the case where $\epsilon=1$, i.e. $\mathrm{MD}(\sigma)=\mathrm{MD}^{1}(\sigma)$, see [15]. The characteristic function of $\mu=\mathrm{MD}^{\epsilon}(\sigma)$ is given by
$\hat{\mu}(z)=\exp\left\{\int_{\mathbb{S}^{d-1}}\int_{0}^{\epsilon}\left(e^{ir\left\langle z,s\right\rangle}-1\right)r^{-1}\mathrm{d}r\sigma(\mathrm{d}s)\right\},\ \ \ z\in\mathbb{R}^{d}.$ (8)
Taking partial derivatives of the exponent in (8) shows that the mean vector and the covariance matrix of $X\sim\mathrm{MD}^{\epsilon}(\sigma)$ are given, respectively, by
$\mathrm{E}[X]=\epsilon\int_{\mathbb{S}^{d-1}}s\sigma(\mathrm{d}s)\ \mbox{ and }\ \mathrm{cov}(X)=\frac{\epsilon^{2}}{2}\int_{\mathbb{S}^{d-1}}ss^{\top}\sigma(\mathrm{d}s).$
One reason for the importance of $\epsilon$-multivariate Dickman distributions is that they satisfy a certain stability property, which stems from the following property of the Lévy measure $D^{\epsilon}$ defined in (7). A simple change of variables shows that for any $\epsilon,\delta>0$ and $B\in\mathfrak{B}(\mathbb{R}^{d})$
$D^{1}(B)=D^{\epsilon}(\epsilon B)\ \mbox{ and }\ D^{\epsilon}(\delta B)=D^{\epsilon/\delta}(B).$
It follows that, for any $\epsilon,\gamma>0$, if $X^{\epsilon}\sim\mathrm{MD}^{\epsilon}(\sigma)$, then
$\frac{1}{\gamma}X^{\epsilon}\sim\mathrm{MD}^{\epsilon/\gamma}(\sigma).$ (9)
Remark 1.
Combining (6) with (9) show that $X^{\epsilon}\sim\mathrm{MD}^{\epsilon}(\sigma)$ if and only if
$X^{\epsilon}\stackrel{{\scriptstyle d}}{{=}}U^{1/\theta}(X^{\epsilon}+\epsilon\xi),$
where $\theta=\sigma(\mathbb{S}^{d-1})$ and $X,\xi,U$ are independent on the right side with $U\sim U(0,1)$ and $\xi\sim\sigma/\theta$. In particular, this means that $\epsilon$-multivariate Dickman distributions belong to the class of multivariate Vervaat perpetuities, see [15].
We now show that $\epsilon$-multivariate Dickman distributions are the only ones that satisfy a scaling property like (9).
Theorem 2.
Let $M\neq 0$ be a Borel measure on $\mathbb{R}^{d}$ such that $M(\{0\})=0$ and for each $h\in(0,\infty)$
$\int_{|x|\leq h}|x|^{2}M(\mathrm{d}x)<\infty.$
Let $M^{\delta}(\mathrm{d}x)=1_{[|x|\leq\delta]}M(\mathrm{d}x)$ for any $\delta>0$. This is a Lévy measure for every $\delta>0$ and we write $X^{\delta}$ to denote a random variable with distribution $\mathrm{ID}(0,M^{\delta},0)$. Assume that there exists a function $\psi:(0,\infty)\mapsto(0,\infty)$ such that for any $\epsilon,\gamma>0$ there exists a non-random $b_{*}(\epsilon,\gamma)\in\mathbb{R}^{d}$ with
$\frac{1}{\psi(\gamma)}X^{\epsilon}\stackrel{{\scriptstyle d}}{{=}}X^{\epsilon/\gamma}+b_{*}(\epsilon,\gamma).$ (10)
Then $\psi(\gamma)=\gamma$ and there exists a spectral measure $\sigma\neq 0$ and a $b_{\epsilon}\in\mathbb{R}^{d}$ such that $X^{\epsilon}+b_{\epsilon}\sim\mathrm{MD}^{\epsilon}(\sigma)$.
Note that a measure $M$ satisfies the assumption of Theorem 2 if and only if $M^{\delta}(\mathrm{d}x)=1_{[|x|\leq\delta]}M(\mathrm{d}x)$ is a Lévy measure for each $\delta>0$. We now show the self-similarity of a process related to $\epsilon$-multivariate Dickman distributions. Let $\sigma$ be a finite Borel measure on $\mathbb{S}^{d-1}$, let $\theta=\sigma(\mathbb{S}^{d-1})$, let $E_{1},E_{2},\dots\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathrm{Exp}(\theta)$, let $\Gamma_{i}=\sum_{k=1}^{i}E_{i}$, and let $\xi_{1},\xi_{2},\dots\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\sigma_{1}:=\sigma/\theta$ be independent of the $E_{i}$’s. In [15] it is shown that
$\sum_{i=1}^{\infty}e^{-\Gamma_{i}}\xi_{i}\sim\mathrm{MD}(\sigma).$ (11)
This can be used to simulate from a multivariate Dickman distribution. However, to do so we must truncate the infinite series. It has been argued, see e.g. [9], that the proper way to truncate is not at a deterministic point. Instead, we should choose some $\tau>0$ and truncate at the random time $N(\tau)$, where $N(\tau)=\max\{i:\Gamma_{i}\leq\tau\}$. This way we have control on the magnitudes of the jumps that we are removing. The remainder in this approximation can be written as
$R_{\tau}=\sum_{i=N(\tau)+1}^{\infty}e^{-\Gamma_{i}}\xi_{i}=\sum_{i=1}^{\infty}1_{[\Gamma_{i}>\tau]}e^{-\Gamma_{i}}\xi_{i}.$
We select $\tau=\tau(t)=\log(1/t)$ for $t\in[0,1]$ and set
$X_{t}=R_{\tau(t)}=\sum_{i=1}^{\infty}1_{[\Gamma_{i}>\log(1/t)]}e^{-\Gamma_{i}}\xi_{i}$ (12)
to be the remainder process. Here we have $X_{0}=0$ a.s. The reason for our choice of $\tau$ is that $\Gamma_{i}>\log(1/t)$ is equivalent to $e^{-\Gamma_{i}}<t$. We now show that $X=\{X_{t}:t\in[0,1]\}$ is a self-similar process.
Theorem 3.
Let $X_{t}$ be as in (12). We have $X_{t}\sim\mathrm{MD}^{t}(\sigma)$ and for any $a\in[0,1]$, we have
$\{X_{at}:t\in[0,1]\}\stackrel{{\scriptstyle d}}{{=}}\{aX_{t}:t\in[0,1]\}.$
4 Approximating Small Jumps of Lévy processes
In which a limit theorem showing that the small jumps of a large class of Lévy processes can be well approximated by a Dickman Lévy process is proved and a related limit theorem for stochastic integral processes is given.

### v2

# On the Simulation of General Multivariate Gamma Distributions using Dickman Approximations

@@ Abstract
We derive a Dickman approximation for the small jumps of a large class of multivariate Lévy processes. We then apply this approximation to develop a simulation method for the class of general multivariate gamma distributions (GMGD). A small-scale simulation study suggests that this method works very well.
Keywords: Multivariate gamma distributions; Dickman distribution; Small jumps of Lévy processes; Simulation
MSC: 60G51; 60F05
@@ 1 Introduction
Gamma distributions and their extensions have been widely used in mathematical finance. Perhaps most prominent is the use of variance gamma and bilateral gamma distributions to model financial returns, see, e.g., [22], [8], [20], [30], and the references therein. In practice, one typically works with a basket of assets and needs a multivariate distribution to jointly model the returns. We focus on the class of general multivariate gamma distributions (GMGD), which is a large class that contains several other models, including the multivariate gamma distributions of [28] as well as important subclasses of the multivariate Thorin class of generalized gamma convolutions (GGC) [3] [21] and of the so-called Class M [23]. GMGD can be used to model baskets of returns directly or it can be first combined with Brownian motion through the process of multivariate subordination, see [4] or [6]. Either way, simulation is an important tool to implement Monte Carlo methods for option pricing and risk estimation. In this paper we develop an approximate simulation method for GMGD, which is based on a multivariate Dickman approximation. The idea is to consider a GMGD Lévy process and to model its small jumps using a multivariate Dickman Lévy process and its large jumps by a compound Poisson process. We will apply this to option pricing in a future work.
For the purposes of simulation, small jumps of Lévy processes are often approximated by simpler processes, see, e.g., Chapter 6 in [8]. The most common process used is Brownian motion, where the approximation is justified by a limit theorem proved in [1] for the univariate case and extended to the multivariate case in [7]. However, this approximation does not hold for gamma and related distributions. In [9] it was shown that, in the univariate case, the small jumps of such Lévy processes can be approximated by Dickman Lévy processes. What has been missing from the literature is a Dickman approximation in the multivariate case. In fact, the multivariate Dickman distribution was only recently introduced in [5] and it was not studied in detail until [17]. In this paper, we prove a limit theorem characterizing when the small jumps of a Lévy process can be approximated by a multivariate Dickman Lévy process, and we show that this always holds for GMGD. Furthermore, we develop a methodology for the simulation of the large jumps in this case, which creates a complete methodology for approximate simulation of GMGD.
The rest of this paper is organized as follows. In Section 2 we review basic facts about Lévy processes and Dickman distributions, and we introduce the slightly more general class of multivariate $\epsilon$-Dickman distributions. In Section 3 we give our main results for approximating the small jumps of certain multivariate Lévy processes by multivariate Dickman Lévy processes and discuss how to use this for simulation. In Section 4 we formally introduce GMGD and give detailed results about simulation using the Dickman approximation. We also perform a small-scale simulation study, which suggests that this method works very well. Proofs are postponed to Section 5.
Before proceeding, we introduce some notation. We write $\mathbb{R}^{d}$ to denote the set of $d$-dimensional column vectors equipped with the usual inner product $\langle\cdot,\cdot\rangle$ and the usual norm $|\cdot|$. We write $\mathbb{S}^{d-1}=\{x\in\mathbb{R}^{d}:|x|=1\}$ to denote the unit sphere in $\mathbb{R}^{d}$, and we write $\mathfrak{B}(\mathbb{R}^{d})$ and $\mathfrak{B}(\mathbb{S}^{d-1})$ to denote the Borel sets on $\mathbb{R}^{d}$ and $\mathbb{S}^{d-1}$, respectively. For a distribution $\mu$ on $\mathbb{R}^{d}$, we write $X\sim\mu$ to denote that $X$ is a random variable with distribution $\mu$ and $X_{1},X_{2},\dots\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mu$ to denote that $X_{1},X_{2},\dots$ are independent and identically distributed (iid) random variables with distribution $\mu$. We write $U(a,b)$ to denote the uniform distribution on $(a,b)$, $\mathrm{Exp}(\lambda)$ to denote the exponential distribution with rate $\lambda$, and $\mathrm{Pois}(\lambda)$ to denote the Poisson distribution with mean $\lambda$. We write $1_{A}$ to denote the indicator function of set $A$ and $\delta_{a}$ to denote the point-mass at $a$. For any $a\in\mathbb{R}$ and $B\subset\mathbb{R}^{d}$ we write $aB=\{ay:y\in B\}$, and for any $C\subset\mathbb{S}^{d-1}$ and $0\leq a<b<\infty$ we write
2 Background
In this sections we review basic facts about Lévy processes and Dickman distributions, and we introduce the slightly more general class of multivariate $\epsilon$-Dickman distributions.
2.1 Lévy Processes
$C_{\mu}(z)=-\langle z,Az\rangle+i\left\langle b,z\right\rangle+\int_{\mathbb{R}^{d}}\left(e^{i\left\langle z,x\right\rangle}-1-i\left\langle z,x\right\rangle 1_{[|x|\leq 1]}\right)M(\mathrm{d}x),\ \ \ z\in\mathbb{R}^{d},$
AA is a $d\times d$-dimensional covariance matrix called the Gaussian part, $b\in\mathbb{R}^{d}$ is the shift, and $M$ is the Lévy measure, which is a Borel measure on $\mathbb{R}^{d}$ satisfying
@@ 2.1 Lévy Processes
The parameters $A$, $M$, and $b$ uniquely determine this distribution and we write $\mu=\mathrm{ID}(A,M,b)$. We call $C_{\mu}$ the cumulant generating function (cgf) of $\mu$. Associated with every infinitely divisible distribution $\mu$ is a Lévy process $\{X_{t}:t\geq 0\}$, where $X_{1}\sim\mu$. In this context, the Lévy measure governs the jumps of the process. Specifically, for any $B\in\mathfrak{B}(\mathbb{R}^{d})$, $M(B)$ is the expected number of jumps that the process has in the time interval $[0,1]$ that fall inside set $B$.
@@ A Lévy process has finite variation if and only if $A=0$ and $M$ satisfies the additional condition
$\int_{\mathbb{R}^{d}}(|x|\wedge 1)M(\mathrm{d}x)<\infty.$
In this case, the cgf can be written in the form
@@
When discussing weak convergence of infinitely divisible distributions and Lévy processes, the concept of vague convergence is fundamental. A portmanteau theorem giving several statements that are equivalent to vague convergence can be found in, e.g., [2]. The definition is as follows.
@@
Let $M_{0},M_{1},M_{2},\dots$ be a sequence of Lévy measures on $\mathbb{R}^{d}$. We say that $M_{n}$ converges vaguely to $M_{0}$ and write $M_{n}\stackrel{{\scriptstyle v}}{{\rightarrow}}M_{0}$ as $n\to\infty$ if $\lim\limits_{n\to\infty}\int_{\mathbb{R}^{d}}f(x)M_{n}(\mathrm{d}x)=\int_{\mathbb{R}^{d}}f(x)M_{0}(\mathrm{d}x)$, for every $f:\mathbb{R}^{d}\mapsto\mathbb{R}$ that is bounded, continuous, and vanishing on a neighborhood of 0.
2.2 Dickman Distributions
@@
For $\epsilon>0$, a random variable $X$ on $\mathbb{R}$ is said to have a generalized $\epsilon$-Dickman distribution if
$X\stackrel{{\scriptstyle d}}{{=}}U^{1/\theta}(X+\epsilon),$
@@
where $\theta\geq 0$ and $U\sim U(0,1)$ is independent of $X$ on the right side. We denote this distribution by $\mathrm{GD}^{\epsilon}(\theta)$. For $\theta=0$, we interpret $\mathrm{GD}^{\epsilon}(0)$ as the distribution concentrated at 0. This is motivated by the fact that, since $U\in(0,1)$ with probability $1$, we have $U^{1/\theta}\to 0$ with probability $1$ as $\theta\downarrow 0$. When $\epsilon=1$, the distribution $\mathrm{GD}^{1}(\theta)$ is called a generalized Dickman distribution and when both $\epsilon=1$ and $\theta=1$, it is just called the Dickman distribution. Many properties and applications for the case $\epsilon=1$ are discussed in the surveys [27], [26], [16], and the references therein. The case $\epsilon\neq 1$ is discussed in [15] and [18]. A multivariate generalization of the Dickman distribution was recently introduced in [5] and it was further studied in [17], where many properties were derived and several approaches for simulation were developed. We now introduce a slight generalization of this model.
@@
For $\epsilon>0$, a random variable $X$ on $\mathbb{R}^{d}$ is said to have a multivariate $\epsilon$-Dickman distribution if
$X\stackrel{{\scriptstyle d}}{{=}}U^{1/\theta}(X+\epsilon\xi),$ (2)
where $\theta\geq 0$ and $X,\xi,U$ are independent on the right side with $U\sim U(0,1)$ and $\xi\sim\sigma_{1}$ for some probability distribution $\sigma_{1}$ on $\mathbb{S}^{d-1}$. Again, for $\theta=0$, we interpret $X$ as having a distribution concentrated on $0\in\mathbb{R}^{d}$, and note that, in this case, the distribution $\sigma_{1}$ does not matter. Let $\sigma=\theta\sigma_{1}$ and note that $\theta=\sigma(\mathbb{S}^{d-1})$ and that for $\theta\neq 0$, $\sigma_{1}=\sigma/\sigma(\mathbb{S}^{d-1})$. Thus, there is no loss of information when working with $\sigma$ instead of $\theta$ and $\sigma_{1}$. We write $\mathrm{MD}^{\epsilon}(\sigma)$ to denote this distribution and we refer to $\sigma$ as the spectral measure. It is readily checked that any nonzero finite Borel measure on $\mathbb{S}^{d-1}$ can serve as the spectral measure of a multivariate $\epsilon$-Dickman distribution. When $\epsilon=1$, $\mathrm{MD}^{1}(\sigma)$ reduces to the multivariate Dickman distribution studied in [5] and [17] and, for simplicity of notation, we write $\mathrm{MD}(\sigma)$ in this case. The generalized $\epsilon$-Dickman distribution $\mathrm{GD}^{\epsilon}(\theta)$ corresponds to $\mathrm{MD}^{\epsilon}(\sigma)$ with dimension $d=1$ and $\sigma=\theta\delta_{1}$. From (2) it is easily checked that $\epsilon$-multivariate Dickman distributions belong to the class of multivariate Vervaat perpetuities, which were introduced in [17].
Lemma 1.
1. For any $\epsilon,\gamma>0$, if $X\sim\mathrm{MD}^{\epsilon}(\sigma)$, then
$\frac{1}{\gamma}X\sim\mathrm{MD}^{\epsilon/\gamma}(\sigma).$ (3)
@@ 1. For any $\epsilon,\gamma>0$, if $X\sim\mathrm{MD}^{\epsilon}(\sigma)$, then
2. If $\mu=\mathrm{MD}^{\epsilon}(\sigma)$, where $\epsilon>0$ and $\sigma$ is a finite Borel measure on $\mathbb{S}^{d-1}$, then $\mu=\mathrm{ID}_{0}(D^{\epsilon},0)$, where
$D^{\epsilon}(B)=\int_{\mathbb{S}^{d-1}}\int_{0}^{\epsilon}1_{B}(rs)r^{-1}\mathrm{d}r\sigma(\mathrm{d}s),\ \ \ B\in\mathfrak{B}(\mathbb{R}^{d}).$ (4)
For $\epsilon=1$, the result in the second part is given in Theorem 5 of [17]. Recall that $D^{\epsilon}(B)$ is the expected number of jumps that the Lévy process associated with $\mathrm{MD}^{\epsilon}(\sigma)$ will have in the time period $[0,1]$, which fall inside of set $B$. Thus, (4) implies that $\epsilon$ is the largest possible magnitude for a jump of this Lévy process.
3 Multivariate Dickman Approximations of Small Jumps